For the function v = 40sin calculate the mean and rms over a range of


1
Expert's answer
2021-05-07T10:24:38-0400

Let function is "v = 40sin\omega t"


Mean value over interval 0 to T where "\omega = \frac{2 \pi}{T}"


"V_{mean} = \frac{\int_0^T vdt}{\int_0^T dt} = \frac{\int_0^T 40sin\omega t dt}{\int_0^T dt} = \frac{\frac{40}{\omega}[-cos\omega t]_0^T}{t_0^T}"

"V_{mean} = \frac{ -40[cos\omega T - cos0]}{\omega T} = -\frac{40}{\omega T}[cos2\pi - cos0] = -\frac{40}{\omega T}[1-1] = 0"



RMS value in interval 0 to T,

"V_{rms} = \sqrt{\frac{\int_0^T v^2 dt}{\int_0^T dt}}"

"V_{rms} = \sqrt{\frac{\int_0^T 1600sin^2\omega t dt}{\int_0^T dt}} = \sqrt{\frac{\int_0^T 1600 \frac{(1-cos2\omega t)}{2}dt}{\int_0^T dt} }"

"V_{rms} = \sqrt{\frac{ 800(t - \frac{1}{2\omega}sin2\omega t)_0^T }{t_0^T}} = \sqrt{\frac{ 800(T - \frac{1}{2\omega}(sin2\omega T - sin0)) }{T} }"

"V_{rms} = \sqrt{\frac{800T}{T}} = \sqrt{800} = 28.284"